OpenStudy (anonymous):

(Find the 15th partial sum of summation of 12i minus 7 from 1 to infinity) PLEASE HELP

OpenStudy (doc.brown):

Sum (add up)\[\sum\]Every 12i-7\[\sum 12i-7\]Starting at i = 1\[\sum _{i=1}12i-7\]for example 12(1)-7 And ending when i = 15\[\sum _{i=1}^{15}12i-7\]for example 12(15)-7. So, \[12(1)-7+12(2)-7+...+12(15)-7\]

OpenStudy (doc.brown):

You don't have to stop at 15, you could keep going to infinity, boring. Stopping at 15 gives you the 15th partial sum.